For moving purposes, the Hendersons bought 22 cardboard boxes for $98.50. There were two kinds of boxes; the large ones cost $5.50 per box, and the small ones cost $3 per box. How many boxes of each kind did they buy?

Answer by Maths68(1474)

Let
Large Box = l
Small Box=s
Total Boxes (large+small)=22
l+s=22
l=22-s………….(1)
cost of large box = 5.50*l
cost of small box=3*s
sum of the cost of (large+small boxes) = total cost
(5.50*l)+(3*s)=98.50
Put the value of l into above equation from (1)
5.50*(22-s)+(3*s)=98.50
121-5.50s+3s=98.50
121-2.5s=98.5
121-98.5=2.5s
22.5=2.5s
s=22.5/2.5
s=9
Put the value of s in (1), we have
l=22-s………….(1)
l=22-9
l=13
13 large boxes and 9 small boxes